OpenStudy (anonymous):

John is using origami paper whose side length is 12 cm. He needs to fold it along the diagonal. How long is the diagonal? John is using origami paper whose side length is 12 cm. He needs to fold it along the diagonal. How long is the diagonal? @Mathematics

OpenStudy (anonymous):

Normally origami papers are square so using the Pythagoras theorem diagonal = sqrt(12^2 + 12^2) = sqrt(144 + 144) = sqrt(288) = 16.97 cm

OpenStudy (anonymous):

diagonal of a square is \[\sqrt{lenght of side ^{2} +lenght of side ^{2}} => \sqrt(144 + 144) =16.9705627 cm ^{2}\]

OpenStudy (anonymous):

12 cm 12 radical 3 cm 12 radical 2 cm 6 cm

OpenStudy (anonymous):

those are the answer choices

OpenStudy (anonymous):

this is the right answer.... 12 radical 2 cm

OpenStudy (anonymous):

\[\sqrt{288} = 12\sqrt{2}\]

OpenStudy (anonymous):

12 radical 2 cm

OpenStudy (anonymous):

thank uu